# A Comparative Analysis of LLM Adaptation: SFT, LoRA, and ICL in Data-Scarce Scenarios

### v1

@@ 3 Experimental Setup and Datasets
Our experiments aim to identify settings where the model successfully learns new information while retaining its general abilities, such as instruction following and performing tasks unrelated to the finetuning data. To explore this systematically, we explore hyperparameters such as the rank for LoRA or learning rate for SFT. We vary the number of training examples provided to the model on a logarithmic scale (log2) to assess the impact of amount of training data on learning and forgetting. For supervised finetuning, we further explore learning rate and training steps, to identify at which a model begins to lose its broader capabilities. We start the exploration with the standard settings (e.g. learning rate 0.001, LoRA rank 4) which is common practice and often set as standard in frameworks. We use Kauldron [^1] https://github.com/google-research/kauldron a popular framework used with Gemma models for our experiments.
We compare the learning paradigms in data-scarce settings which crucially, permits a head-to-head comparison with in-context learning (ICL) and use up to 32k tokens with the Gemma-3 models to accommodate the same number of examples. For a controlled comparison, we therefore run SFT and LoRA with an identical example budget and then extend the analysis to larger datasets to characterize scaling beyond this common baseline.
@@ 5 Discussion
Hypothesis and insights. High-level, task-specific decision rules live in late layers (Tenney et al., 2019). Low-rank updates inject changes where they most directly affect the logits, and these later layers are closest to the output. With cross-entropy loss, gradients diminish with depth during backpropagation. RMSNorm further attenuates them in early blocks, while residual connections might mitigate this in Gemma. Conversely, early layers encode reusable lexical and orthographic features that are already sufficient for the task, meaning changing them offers little gain. Attention vs. MLP roles differ; attention layers adjust routing/aggregation while MLPs adjust feature re-weighting.

### v2

@@ 3 Experimental Setup and Datasets
Our experiments aim to identify settings where the model successfully learns new information while retaining its general abilities, such as instruction following and performing tasks unrelated to the finetuning data. To explore this systematically, we explore hyperparameters such as the rank for LoRA or learning rate for SFT. We vary the number of training examples provided to the model on a logarithmic scale (log2) to assess the impact of the amount of training data on learning and forgetting. For supervised finetuning, we further explore learning rate and training steps, to identify the point at which a model begins to lose its broader capabilities. We start the exploration with the standard settings (e.g. learning rate 0.001, LoRA rank 4) which is common practice and often set as standard in frameworks. We use Kauldron [^1] https://github.com/google-research/kauldron a popular framework used with Gemma models for our experiments.
We compare the learning paradigms in data-scarce settings which crucially, permits a head-to-head comparison with in-context learning (ICL) and use up to 32k tokens with the Gemma-3 models to accommodate the same number of examples. We use Gemma 4B model throughout the paper. For a controlled comparison, we therefore run SFT and LoRA with an identical example budget and then extend the analysis to larger datasets to characterize scaling beyond this common baseline.
@@ 5 Discussion
Figure 13: Evolution of SFT weight update ($\Delta W$) magnitudes for attention (top row) and MLP (bottom row) weights. Each heatmap plots update magnitude across model layers (y-axis) against training samples (x-axis), with plots from left to right showing training at 800, 1600, and 3000 steps.
Hypothesis and insights. High-level, task-specific decision rules live in late layers (Tenney et al., 2019). Low-rank updates inject changes where they most directly affect the logits, and these later layers are closest to the output. With cross-entropy loss, gradients diminish with depth during backpropagation. RMSNorm further attenuates them in early blocks, while residual connections might mitigate this in Gemma. Conversely, early layers encode reusable lexical and orthographic features that are already sufficient for the task, meaning changing them offers little gain. Attention vs. MLP roles differ; attention layers adjust routing/aggregation while MLPs adjust feature re-weighting. This version has the same core issues as the previous text (typos, run-on sentences). Here is a corrected version and a breakdown of the changes. Figure 13 shows the evolution of the weight updates for SFT and allows for comparison with LoRA. These changes are substantially larger than for LoRA. For comparison, the scale on the right side of the y-axis for LoRA is at most 0.3 for the L1-Norm; in contrast, the changes by SFT are more than an order of magnitude higher. Similar to LoRA, the majority of the changes occur in the attention weights.